### v2

@@ Abstract
We present two neural network approaches that approximate the solutions of static and dynamic conditional optimal transport (COT) problems. Both approaches enable conditional sampling and conditional density estimation, which are core tasks in Bayesian inference—particularly in the simulation-based (“likelihood-free”) setting. Our methods represent the target conditional distributions as transformations of a tractable reference distribution and, therefore, fall into the framework of measure transport. Although many measure transport approaches model the transformation as COT maps, obtaining the map is computationally challenging, even in moderate dimensions. To improve scalability, our numerical algorithms use neural networks to parameterize COT maps and further exploit the structure of the COT problem. Our static approach approximates the map as the gradient of a partially input-convex neural network. It uses a novel numerical implementation to increase computational efficiency compared to state-of-the-art alternatives. Our dynamic approach approximates the conditional optimal transport via the flow map of a regularized neural ODE; compared to the static approach, it is slower to train but offers more modeling choices and can lead to faster sampling. We demonstrate both algorithms numerically, comparing them with competing state-of-the-art approaches, using benchmark datasets and simulation-based Bayesian inverse problems.
@@ 1 Introduction
Most methods for sampling the posterior, such as Markov chain Monte Carlo (MCMC) [7], or for otherwise approximating it in some variational family [13], require evaluations of the unnormalized posterior density. This requirement limits such methods to the “conventional” setting of a tractable likelihood function and prior density—i.e., a likelihood and prior that can be evaluated pointwise. This setting includes likelihoods common in inverse problems, given by a forward operator that maps ${\bf x}$ to a prediction of the observations, which is then perturbed by noise. Even in this setting, producing thousands of approximately i.i.d. samples from the posterior distribution may require millions of unnormalized posterior density evaluations. This can be especially expensive for science and engineering applications where the forward operator entering the likelihood involves the solution of differential equations.
@@
Conditional sampling is even more challenging when the likelihood model—and hence the unnormalized posterior density—is intractable to compute. One common reason is that the forward model is intrinsically stochastic, or (equivalently) that it contains perhaps a high-dimensional set of random latent parameters ${\bf c}$ that are not the target of inference; in this setting, computing the likelihood ${\bf x}\mapsto\pi({\bf y}|{\bf x})$ requires an expensive step of marginalization. In other applications, often the case in data assimilation [53] or inverse problems with data-driven priors [23], the prior distribution may be known only through samples with no closed-form density. In both of these cases, one can simulate sample pairs $({\bf x}_{i},{\bf y}_{i})\sim\pi({\bf x},{\bf y})$ but not evaluate their joint density.
@@
As a possible remedy, likelihood-free (also called simulation-based) inference methods [12] have gained significant recent interest, Instead of relying on evaluations of the unnormalized posterior density, most of these methods employ some form of sample-driven measure transport, either as a way of estimating conditional densities [43] or directly sampling from the posterior [18]. The latter can understood as using transport for conditional generative modeling: we seek transport maps, amortized over multiple realizations of the data the data ${\bf y}$, that push forward a given reference distribution to the target posterior distribution $\pi({\bf x}|{\bf y})$. Once such a map is obtained, one can readily generate i.i.d. samples from the target by evaluating the map on samples drawn from the reference given any realization of data ${\bf y}$. This distinguishes transport maps from methods such as MCMC, which are intrinsically serial (thus challenging to parallelize) and typically requiring re-initialization for any new value of ${\bf y}$.
@@
Under mild assumptions on the reference and posterior, there exist infinitely many transport maps that couple the two distributions. Accordingly, many parameterizations and construction methods for such transport maps have been pursued in recent literature. In particular, neural networks are widely used to represent transport maps, due to their expressiveness and scalability to high dimensions [2, 20, 14, 40, 24]. Yet in many settings, the target map, e.g., the map obtained in the limit of increasing expressiveness of the network, is not clearly or uniquely specified. The lack of a specific target transport leaves open a wide variety of ad hoc design choices, where these choices implicitly determine the class of transport maps being considered. Moreover, it is then challenging to understand the approximation properties of resulting transports and their implications for computational cost. A more systematic way of using neural networks to represent conditional transports with a well-specified target map is desirable. To this end, we will seek neural network approximations of conditional optimal transport (COT) maps, in particular, maps that are optimal under the $L^{2}$ transport cost. These maps are unique, under mild assumptions on the reference and the target conditionals, and are also known as conditional Brenier maps [9].
@@
This paper contributes two neural network approaches for learning the static and dynamic COT maps, respectively, through maximum likelihood training. We select conditional Brenier maps as our target transport class and enforce their gradient-of-convex-potential structure to regularize the transport map learning problem, eliminating exploration of the unnecessary off-target choices. This leads directly to our first approach, the partially convex potential map (PCP-Map), where we parameterize the map as the (partial) gradient of a partially input convex neural network (PICNN) [1]. PCP-Maps, by construction, approximate static conditional Brenier maps in the limit of network expressiveness. For our second approach, called conditional optimal transport flow (COT-Flow), we extend our prior approach, OT-Flow [40], to include amortization over multiple realizations of the data. Like OT-Flow, our method considers dynamic measure transport, which has recently gained much popularity. The dynamic approach realizes transport by integrating a velocity field for a finite time. COT-Flow seeks a conditional optimal transport defined by the flow map of an ODE with constant-velocity (in a Lagrangian sense) trajectories. Rather than strongly enforcing the structure of optimal transport on this velocity field, we encourage optimality of the conditional transport via appropriate penalties on the velocity. In particular, we represent the velocity as the gradient of a scalar potential parameterized by a neural network (hence obtaining a neural ODE) and penalize the squared magnitude of the velocity and violations of a Hamilton–Jacobi–Bellman PDE.
@@
We compare the performance of our approaches with each other and against other state-of-the-art neural and non-neural approaches for conditional sampling and density estimation in terms of the computational cost of training and sampling, the quality of the resulting samples, and the complexity of hyperparameter tuning. We conduct numerical experiments using six UCI tabular datasets [32], inverse problems involving the stochastic Lotka-Volterra model [44], and the one-dimensional shallow water equations [48]. Across all experiments, we observe that both PCP-Map and COT-Flow achieve competitive accuracy as measured by negative log-likelihoods, sample visualizations, and posterior calibration analyses, compared to consistent approximate Bayesian computation methods [57] and other transport or deep learning approaches [3, 48]. While PCP-Map and COT-Flow enjoy comparable accuracy, the strongly-constrained PCP-Map can be trained more quickly than COT-Flow in our experiments. Also, PCP-Map relies on fewer and less influential hyperparameters than the weakly-constrained COT-Flow and thus requires less parameter tuning. COT-Flow, however, offers faster sampling after training due to its nearly straight-line velocity field. We also conduct a comparative study of PCP-Map and a conditional version of the optimal transport approach proposed in [24] and observe that our numerical implementation improves the numerical accuracy and efficiency.
@@ 2 Background and related work
With this construction, we can sample from the conditional distribution $\pi({\bf x}|{\bf y})$ simply by evaluating the mapping $g:{\bf z}_{x}\mapsto\mathsf{g}({\bf z}_{x},{\bf y})$ for ${\bf z}_{x}\sim\rho_{{\bf z}_{x}}({\bf z}_{x})$. We will focus on such a $g$ in this paper, and call it the conditional generator (interchangeably with ‘conditional transport map’). We view $g$ as a mapping from $\mathbb{R}^{n}\to\mathbb{R}^{n}$, parameterized by ${\bf y}$, and will write $g({\bf z}_{x};{\bf y})$, for ${\bf z}_{x}\in\mathbb{R}^{n}$, to make this interpretation clear. We will also use $g^{-1}(\cdot\,;{\bf y})$ to denote the inverse of this mapping (when it exists) for any given value of ${\bf y}$. In other words, $g\left(g^{-1}({\bf x};{\bf y});{\bf y}\right)={\bf x}$.
@@ 3 Partially convex potential maps (PCP-Map)
Our first approach, PCP-Map, approximates the static COT problem by solving a maximum likelihood estimation problem over a class of neural network-based conditional transport maps whose structure is enforced to match the structure of conditional Brenier maps—which are partial gradients of partially convex potentials. By the neural network architectural design, this partial convexity is guaranteed for any choice of the network weights.
@@
Given samples from the joint distribution, our algorithm seeks a conditional generator $g$ by solving the maximum likelihood problem
(5) $\min_{g^{-1}\in\mathcal{G}}\;J_{{\rm NLL}}[g],$
@@
(6) $J_{{\rm NLL}}[g]=\mathbb{E}_{\pi({\bf x},{\bf y})}\left[{\frac{1}{2}}\left\|g^{-1}({\bf x};{\bf y})\right\|^{2}-\log\det\nabla_{\bf x}g^{-1}({\bf x};{\bf y})\right].$
This objective arises from choosing $\rho_{{\bf z}_{x}}$ to be a standard Gaussian and it maximizes the likelihood that $g^{-1}$ samples are standard Gaussian. Since the learning problem in (6) only involves the inverse conditional generator $g^{-1}$, in (5) we follow [2] and directly seek conditional generators whose inverses take the structural form of conditional Brenier maps, i.e.,
(7) $\mathcal{G}=\left\{({\bf x};{\bf y})\mapsto\nabla_{{\bf x}}G({\bf x};{\bf y})\ \text{s.t.}\ G:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}\text{is convex in its first argument}\right\}.$
These parameterizations are equivalent: if, for any fixed ${\bf y}$, $g(\cdot;{\bf y})$ is the gradient of a convex potential, then $g(\cdot;{\bf y})^{-1}$ is also the gradient of a (different) convex potential by Brenier’s theorem. While this choice avoids inverting the generator during training, sampling requires inverting the learned map. Note that the inverse conditional generator pushes forward the conditional distribution $\pi({\bf x}|{\bf y})$ to the reference distribution. We note that the negative log-likelihood $J_{\rm NLL}$ also agrees (up to an additive constant) with the Kullback–Leibler (KL) divergence from the pushforward $g(\cdot;{\bf y})_{\sharp}\rho_{{\bf z}_{x}}$ to the conditional distribution $\pi({\bf x}|{\bf y})$, in expectation over the conditioning variable ${\bf y}\sim\pi({\bf y})$.
@@
In this work, we directly enforce the structural form of conditional Brenier maps on the inverse conditional generator using scalar-valued partially input convex neural networks (PICNNs) [1]. In particular, we parameterize $g^{-1}_{\theta}$ as the gradient of a PICNN based potential $\tilde{G}_{\theta}:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ that depends on weights $\theta$. A PICNN is a feed-forward neural network that is specifically designed to ensure convexity in some of its inputs; see the original work that introduced this neural network architecture in [1] and its use for generative modeling in [24]. To the best of our knowledge, investigating if PICNNs are universal approximators of partially input convex functions is still an open issue, also see [28, Section IV], that is beyond the scope of our paper; a perhaps related result for fully input convex neural networks is given in [24, Appendix C].
To ensure the monotonicity of $g^{-1}_{\theta}$, we construct $\tilde{G}_{\theta}$ to be strictly convex as a linear combination of a scalar-valued PICNN and a positive definite quadratic term. That is,
@@
Table 1: Parameter dimensions of a $K$-layer PICNN architecture from (10). Since the dimensions of the inputs, ${\bf x}\in\mathbb{R}^{n}$ and ${\bf y}\in\mathbb{R}^{m}$, are given, and the network outputs a scalar, we vary only the depth, $K$, feature width, $w$, and context width, $u$, in our experiments.
@@
To compute the log determinant of $g^{-1}_{\theta}$, we use vectorized automatic differentiation to obtain the Hessian of $\tilde{G}_{\theta}$ with respect to its first input and then compute its eigenvalues. This is feasible when the Hessian is moderate; e.g., in our experiments, it is less than one hundred. We use efficient implementations of these methods that parallelize the computations over all the samples in a batch.
@@
Due to our neural network parameterization, there is generally no closed-form relation between $g_{\theta}$ and $g^{-1}_{\theta}$. As in [24], we approximate the inverse of $g^{-1}_{\theta}$ during sampling as the Legendre-Fenchel dual. That is, given a reference sample ${\bf z}_{x}\sim\rho_{{\bf z}_{x}}({\bf z}_{x})$ and a conditioning variable ${\bf y}\sim\pi({\bf y})$, we solve the convex optimization problem
(12) ${\bf v}^{*}=\operatorname*{arg\,min}_{{\bf v}}\;\tilde{G}_{\theta}({\bf v};{\bf y})-{\bf z}_{x}^{\top}{\bf v}.$
@@
In our numerical experiments, we vary only three hyperparameters to adjust the complexity of the architecture. As described in Section 5 we randomly sample the depth, $K$, the feature width, $w$, and the context width, $u$, from the values in Table 2.

### v3

@@ Abstract
We present two neural network approaches that approximate the solutions of static and dynamic conditional optimal transport (COT) problems. Both approaches enable conditional sampling and conditional density estimation, which are core tasks in Bayesian inference—particularly in the simulation-based (“likelihood-free”) setting. Our methods represent the target conditional distribution as a transformation of a tractable reference distribution. Obtaining such a transformation, chosen here to be an approximation of the COT map, is computationally challenging even in moderate dimensions. To improve scalability, our numerical algorithms use neural networks to parameterize candidate maps and further exploit the structure of the COT problem. Our static approach approximates the map as the gradient of a partially input-convex neural network. It uses a novel numerical implementation to increase computational efficiency compared to state-of-the-art alternatives. Our dynamic approach approximates the conditional optimal transport via the flow map of a regularized neural ODE; compared to the static approach, it is slower to train but offers more modeling choices and can lead to faster sampling. We demonstrate both algorithms numerically, comparing them with competing state-of-the-art approaches, using benchmark datasets and simulation-based Bayesian inverse problems.
@@ 1 Introduction
Most methods for sampling the posterior, such as Markov chain Monte Carlo (MCMC) [7], or for otherwise approximating it in some variational family (e.g., [13]), require evaluations of the unnormalized posterior density. This requirement limits such methods to the “conventional” setting of a tractable likelihood function and prior density—i.e., a likelihood and prior that can be evaluated pointwise. This setting includes likelihoods common in inverse problems, given by a forward operator that maps ${\bf x}$ to a prediction of the observations, which is then perturbed by noise. Even in this setting, producing thousands of approximately i.i.d. samples from the posterior distribution may require millions of unnormalized posterior density evaluations. This can be especially expensive for science and engineering applications where the forward operator entering the likelihood involves solving differential equations.
@@
Conditional sampling is even more challenging when the likelihood—and hence the unnormalized posterior density—is intractable to compute. One common reason is that the forward model is intrinsically stochastic or (equivalently) that it contains a set of (possibly high-dimensional) random latent parameters ${\bf c}$ that are not the target of inference; in this setting, computing the likelihood ${\bf x}\mapsto\pi({\bf y}|{\bf x})$ requires an expensive step of marginalization. In other applications, often the case in data assimilation [56] or inverse problems with data-driven priors [24], the prior distribution may be known only through samples, without a closed-form density. In both of these cases, one can simulate sample pairs $({\bf x}_{i},{\bf y}_{i})\sim\pi({\bf x},{\bf y})$ but not evaluate their joint density.
@@
As a possible remedy, likelihood-free (also called simulation-based) inference methods [12] have seen significant recent interest. Instead of relying on evaluations of the unnormalized posterior density, most of these methods employ some form of sample-driven measure transport, either as a way of estimating conditional densities [44] or directly sampling from the posterior [18]. The latter can be understood as using transport for conditional generative modeling: we seek transport maps, amortized over multiple realizations of the data ${\bf y}$, that push forward a given reference distribution to the target posterior distribution $\pi({\bf x}|{\bf y})$. Once such a map is obtained, one can readily generate i.i.d. samples from the target by evaluating the map on samples drawn from the reference given any realization of data ${\bf y}$. This distinguishes transport maps from methods such as MCMC, which are intrinsically serial (thus challenging to parallelize) and typically require re-initialization for any new value of ${\bf y}$.
@@
Under mild assumptions on the reference and posterior distributions, there exist infinitely many transport maps that couple the two. Accordingly, many parameterizations and construction methods for such transport maps have been pursued in recent literature. In particular, neural networks are widely used to represent transport maps, due to their expressiveness and scalability to high dimensions [2, 20, 14, 41, 25]. Yet in many settings, the target map, e.g., the map obtained in the limit of increasing expressiveness of the network, is not clearly or uniquely specified. The lack of a specific target transport leaves open a wide variety of ad hoc design choices, where these choices implicitly determine the class of transport maps being considered. Moreover, it is challenging to understand the approximation properties of resulting transports and their implications for computational cost. A more systematic way of using neural networks to represent conditional transports with a well-specified target map is desirable. To this end, we will seek neural network approximations of conditional optimal transport (COT) maps, in particular, maps that are optimal under the $L^{2}$ transport cost. These maps are unique, under mild assumptions on the reference and the target conditionals, and are also known as conditional Brenier maps [9].
@@
This paper contributes two neural network approaches for learning the static and dynamic COT maps, respectively, through maximum likelihood training. We select conditional Brenier maps as our target transport class and enforce their gradient-of-convex-potential structure to regularize the transport map learning problem, eliminating exploration of the unnecessary off-target choices. This leads directly to our first approach, the partially convex potential map (PCP-Map), where we parameterize the map as the (partial) gradient of a partially input convex neural network (PICNN) [1]. PCP-Maps, by construction, approximate static conditional Brenier maps in the limit of network expressiveness. For our second approach, called conditional optimal transport flow (COT-Flow), we extend our previous approach, OT-Flow [41], to include amortization over multiple realizations of the data. Like OT-Flow, our method considers dynamic measure transport, which has recently gained much popularity. The dynamic approach realizes transport by integrating a velocity field for a finite time. COT-Flow seeks a conditional optimal transport defined by the flow map of an ODE with constant-velocity (in a Lagrangian sense) trajectories. Rather than strongly enforcing the structure of optimal transport on this velocity field, we encourage optimality of the conditional transport via appropriate penalties on the velocity. In particular, we represent the velocity as the gradient of a scalar potential parameterized by a neural network (hence obtaining a neural ODE) and penalize the squared magnitude of the velocity and violations of a Hamilton–Jacobi–Bellman PDE.
@@
We compare the performance of our approaches with each other and against other state-of-the-art neural and non-neural approaches for conditional sampling and density estimation in terms of the computational cost of training and sampling, the quality of the resulting samples, and the complexity of hyperparameter tuning. We conduct numerical experiments using six UCI tabular datasets [28], inverse problems involving the stochastic Lotka–Volterra model [45], and the one-dimensional shallow water equations [50]. Across all experiments, we observe that both PCP-Map and COT-Flow achieve competitive accuracy as measured by negative log-likelihoods, sample visualizations, and posterior calibration analyses, compared to consistent approximate Bayesian computation methods [60] and other transport or deep learning approaches [3, 50]. While PCP-Map and COT-Flow enjoy comparable accuracy, the strongly-constrained PCP-Map can be trained more quickly than COT-Flow in our experiments. Also, PCP-Map relies on fewer and less influential hyperparameters than the weakly constrained COT-Flow and thus requires less hyperparameter tuning. COT-Flow, however, offers faster sampling after training due to its nearly straight-line velocity field. We also conduct a comparative study of PCP-Map and a conditional version of the optimal transport approach proposed in [25] and observe that our numerical implementation improves numerical accuracy and efficiency.
@@ 2 Background and related work
With this construction, we can sample from the conditional distribution $\pi({\bf x}|{\bf y})$ simply by evaluating the mapping $g:{\bf z}_{x}\mapsto\mathsf{g}({\bf z}_{x},{\bf y})$ for ${\bf z}_{x}\sim\rho_{{\bf z}_{x}}({\bf z}_{x})$. We will focus on such a $g$ in this paper, and call it the conditional generator. We view $g$ as a mapping from $\mathbb{R}^{n}\to\mathbb{R}^{n}$, parameterized by ${\bf y}$, and will write $g({\bf z}_{x};{\bf y})$, for ${\bf z}_{x}\in\mathbb{R}^{n}$, to make this interpretation clear. We will also use $g^{-1}(\cdot\,;{\bf y})$ to denote the inverse of this mapping (when it exists) for any given value of ${\bf y}$. In other words, $g\left(g^{-1}({\bf x};{\bf y});{\bf y}\right)={\bf x}$.
@@ 3 Partially convex potential maps (PCP-Map)
Our first approach, PCP-Map, approximates the conditional optimal transport map as the partial gradient of a partially input convex neural network (PICNN) [1]. This design enforces the structure of conditional Brenier maps. As common in normalizing flows, we learn the parameters of the neural network by maximizing the likelihood of the training samples.
@@
Given samples from the joint distribution, we train the conditional generator $g$ by solving the maximum likelihood problem
(5) $\min_{g^{-1}\in\mathcal{G}}\;J_{{\rm NLL}}[g^{-1}],$
@@
(6) $J_{{\rm NLL}}[g^{-1}]=\mathbb{E}_{\pi({\bf x},{\bf y})}\left[{\frac{1}{2}}\left\|g^{-1}({\bf x};{\bf y})\right\|^{2}-\log\det\nabla_{\bf x}g^{-1}({\bf x};{\bf y})\right].$
This objective arises from choosing $\rho_{{\bf z}_{x}}$ to be a standard Gaussian; maximizing it seeks to make $g^{-1}({\bf x};{\bf y})$ also standard Gaussian. Since the learning problem in (6) only involves the inverse conditional generator $g^{-1}$, in (5) we follow [2] and directly seek conditional generators whose inverses have the structure of conditional Brenier maps, i.e.,
(7) $\mathcal{G}=\left\{{\bf x}\mapsto\nabla_{{\bf x}}G({\bf x},{\bf y})\ \text{s.t.}\ G:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}\text{ is convex in ${\bf x}$}\right\}.$
These parameterizations are equivalent: if, for any fixed ${\bf y}$, $g(\cdot;{\bf y})$ is the gradient of a convex potential, then $g(\cdot;{\bf y})^{-1}$ is also the gradient of a (different) convex potential by Brenier’s theorem, and vice versa. While this choice avoids inverting the generator during training, sampling requires inverting the learned map. Note that the inverse conditional generator pushes forward the conditional distribution $\pi({\bf x}|{\bf y})$ to the reference distribution. We note that the negative log-likelihood $J_{\rm NLL}$ also agrees (up to an additive constant) with the Kullback–Leibler (KL) divergence from the pushforward $g(\cdot;{\bf y})_{\sharp}\rho_{{\bf z}_{x}}$ to the conditional distribution $\pi({\bf x}|{\bf y})$, in expectation over the conditioning variable ${\bf y}\sim\pi({\bf y})$.
@@
We directly parameterize $g^{-1}_{\theta}$ as the partial gradient of a scalar-valued PICNN-based potential $\tilde{G}_{\theta}:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ that depends on weights $\theta$. A PICNN is a feed-forward neural network that is specifically designed to ensure convexity in part of its inputs; see the original work that introduced this neural network architecture in [1] and its use for generative modeling in [25]. To the best of our knowledge, investigating whether PICNNs are universal approximators of partially input convex functions remains an open issue (also see [30, Section IV]) that is beyond the scope of our paper; a perhaps related result for fully input convex neural networks is given in [25, Appendix C].
To ensure the strict monotonicity of $g^{-1}_{\theta}$, we construct $\tilde{G}_{\theta}$ to be strictly convex as a linear combination of a scalar-valued PICNN and a positive definite quadratic term. That is,
@@
Table 1: PCP-Map, Section 3. Parameter dimensions of a $K$-layer PICNN architecture from (10). Since the dimensions of the inputs, ${\bf x}\in\mathbb{R}^{n}$ and ${\bf y}\in\mathbb{R}^{m}$, are given, and the network outputs a scalar, we vary only the depth, $K$, feature width, $w$, and context width, $u$, in our experiments.
@@
To compute the log determinant of $g^{-1}_{\theta}$, we use vectorized automatic differentiation to obtain the Hessian of $\tilde{G}_{\theta}$ with respect to its first input and then compute its eigenvalues. This is feasible when the Hessian’s dimension is moderate; e.g., in our experiments, it is less than one hundred. We use efficient implementations of these methods that parallelize the computations over all the samples in a batch.
@@
Due to the neural network parameterization, there is generally no closed-form relation between $g_{\theta}$ and $g^{-1}_{\theta}$. As in [25], we approximate the inverse of $g^{-1}_{\theta}$ during sampling as the Legendre-Fenchel dual. That is, given a reference sample ${\bf z}_{x}\sim\rho_{{\bf z}_{x}}({\bf z}_{x})$ and a conditioning variable ${\bf y}\sim\pi({\bf y})$, we generate samples by solve the convex optimization problem
(12) ${\bf v}^{*}=\operatorname*{arg\,min}_{{\bf v}}\;\tilde{G}_{\theta}({\bf v},{\bf y})-{\bf z}_{x}^{\top}{\bf v}.$
@@
Numerically, the optimization problem (12) is solved approximately using $l$-BFGS with line search under strong Wolfe conditions. The optimization is terminated when a pre-set tolerance is reached.
@@
In our numerical experiments, we vary only three hyperparameters to adjust the complexity of the architecture. As described in Section 5 we select the depth $K$, feature width $w$, and context width $u$ from the values in Table 2.